OpenStudy (usukidoll):

All I know is that x = 1 is a root of the polynomial, but I don't know what to do next. Consider the sequence of integers $$ a_{n} $$ for $$ n \geq 0 $$ defined recursively as follows:\\ $$ a_{0}=a_{1}=a_{2}=0; a_{3}=1. $$\\ For $$ n \geq 4, $$\\ $$ a_{n} = 6a_{n-1}-11a_{n-2}+6a_{n-3} $$\\ Find a closed formula for $$ a_{n} $$ (Note that x = 1 is a root of the polynomial $$ 6x^3-11x^2+6x-1.) $$\\

OpenStudy (kainui):

Latex broken? Can you repost it as a comment

OpenStudy (usukidoll):

efff latex is dead in here -_-

OpenStudy (usukidoll):

Consider the sequence of integers $$ a_{n} $$ for $$ n \geq 0 $$ defined recursively as follows:\\ $$ a_{0}=a_{1}=a_{2}=0; a_{3}=1. $$\\ For $$ n \geq 4, $$\\ $$ a_{n} = 6a_{n-1}-11a_{n-2}+6a_{n-3} $$\\ Find a closed formula for $$ a_{n} $$ (Note that x = 1 is a root of the polynomial $$ 6x^3-11x^2+6x-1.) $$\\

OpenStudy (usukidoll):

either I have amnesia or I really don't know how this works... but I know that x = 1 is a root

OpenStudy (kainui):

Oh you could solve this with generating functions it looks like that's how they're wanting you to solve it maybe. At least you can factor out (x-1) from that polynomial maybe to turn it into like partial fraction decomposition if that makes sense to you. I'm gone for mothers day stuff now, good luck in the mean time!

OpenStudy (usukidoll):

like finding all roots to that polynomial?

OpenStudy (kainui):

Ok so things are slow to start right now, so we're good, define this sequence: \[S(x) = \sum_{n=0}^\infty a_n x^n\]\[S(x) = a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+\cdots\] Now if you add this thing to itself, but shifted and multiplied, you'll end up with the recurrence relation. \[S(x) = \text{some extra terms}+ 6xS(x)-11x^2S(x)+6x^3S(x)\] The extra terms come out of shifting the polynomials over to get them to cancel out, then you can factor and use partial fraction decomposition to get a simpler closed form probably. I gotta go for real this time though so good luck haha

OpenStudy (usukidoll):

:S so that can be used to find the remaining roots.. maybe I should search recurrence relation first haha (3x-1)(2x-1)(x-1) x = 1, 1/2, 1/3

OpenStudy (usukidoll):

:S!

OpenStudy (kainui):

OK so did you get anywhere or are you still sorta flopping around on this

OpenStudy (kainui):

Ok to be fair maybe that's a loaded question hahaha I wouldn't've known what to do if I didn't have someone show me this trick either. I learned it in this book called 'generatingfunctionology' so let's get to it.

OpenStudy (kainui):

The idea is this, start here by defining this object for ourselves: \[S(x) = a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+\cdots\] Now what we do is we play around with this by some tricks, so that by adding this thing to itself we end up getting it to simplify due to the recursion relation we were given, \[a_n = 6a_{n-1} - 11 a_{n-2}+6a_{n-3}\] So as a specific example, we'd want when n=4 \[a_4=6a_3-11a_2+6a_1\] It's all about writing the sum the right way so that it matches up to follow a pattern. So try to figure out, what do you have to multiply S(x) by to make it so that you get those terms to line up like that? The idea is by multiplying by a power of x, you shift the terms around. For instance side example: \[T(x) = b_0+b_1x+b_2x^2+\cdots\] \[2xT(x)=2b_0x+2b_1x^2+2b_2x^3+\cdots\] combined gives: \[T(x)+2xT(x) = b_0 + (b_1+2b_0)x+(b_2+2b_1)x^2+\cdots\]

OpenStudy (usukidoll):

are we getting a telescoping series out of this?

OpenStudy (kainui):

no, we're getting a recurrence relation outta it. Look at each coefficient as being a specific version of the recurrence relation you are looking for in general

OpenStudy (usukidoll):

so if n = 4 what our result should be is \[a_{4} =6a_{3}-11a_{2}+6a_{1}\] we are given that a_0=a_1=a_2=0 and a_3 = 1 so a_4 should be 6 by itself. ?!

OpenStudy (kainui):

Well, relax for a second on that, we'll sorta have to treat the initial terms separately but we'll come to that

OpenStudy (kainui):

Pretend the initial stuff doesn't matter for now, all I want you to do is come up with a way write it so that the terms add up the right way. Find the things to multiply S(x) by so that you get the recurrence relation you need

OpenStudy (usukidoll):

I could have one as x and then another as 5x but adding that together will be 6x and that would only work for the first and third terms sinc e both are 6